# Efficient and Intuitive Two-Phase Validation Across Multiple Models via Principal Components

### v1

@@ 2.3 Imputation Procedure
In our validated subset, we can estimate the $j$th linear regression model in (1) using the observed exposure $X_{ji}$ directly. However, for the rest of the study, we will instead replace their missing $X_{ji}$ value with an informative placeholder. Specifically, we adopt a single deterministic imputation framework in the Simulations (Section 3), wherein unvalidated patients contribute predicted exposures $\widehat{X}_{ji}$ to the analysis models. (Single imputation is computationally quicker than multiple, and we focused on the estimates’ empirical variability.) However, multiple imputation is preferred if estimated variances are desired (e.g., for inference), in which case the imputation models would need to be modified. R code for imputation under each design can be found online at https://github.com/sarahlotspeich/ETS_PCA.
@@ 2.3.1 Different Imputation Models Under Different Validation Study Designs
For an unvalidated patient, $\widehat{X}_{ji}$ comes from an imputation model fit to the validation study. Any fully available information that helps estimate $X_{ji}$ can be incorporated, and any variables used in the validation study design must be included [zhou2002semiparametric].
Under an SRS design, the imputed values are defined simply from the error-prone exposure $X_{ji}^{*}$ and the additional confounders ${\mathbf{Z}}_{i}$ (presumed to be error-free and fully observed) as
@@
When $X_{ji}$ is numeric, the estimated coefficients in (2) can be obtained via normal linear regression using only data from the validation subset. Of course, different regression models could instead be used to impute other types of exposures, like binary or categorical.
If the validation study was selected under an ETS-$X_{j}^{*}$ design, the error-prone exposure $X_{j}^{*}$ used in the design must be included in all exposures’ imputation models. Let $j^{\prime}$ index which model was used in the design ($j^{\prime}\in\{1,\dots,P\}$). Then, the imputed values $\widehat{X}_{ji}$ are:
@@
Note that the coefficients will take on different values in the first versus second case in (3).
@@
$\widehat{X}_{ji}$ $=\widehat{\alpha}_{0j}+\widehat{\alpha}_{1j}X^{*}_{ji}+\widehat{\alpha}_{2j}Z_{i}+\widehat{\alpha}_{3j}PC^{*}_{i1}.$ (4)
As with the SRS design, estimating the imputation models under either ETS design depends on the type of covariates, but can be done using standard regression techniques. For ease of notation, $\widehat{\boldsymbol{\alpha}}$ are the estimated coefficients for any of the imputation models in (2)–(4).
Single imputation following (2)–(4) has been called deterministic imputation, since missing values are replaced with a single model prediction $\widehat{X}_{ji}$ and treated as fixed in the subsequent analyses [DAgostinoMcGowan2024]. This approach can be implemented using existing software to (i) fit the imputation model and (ii) calculate the predictions. Alternatively, the mice function in the mice R package can be used with argument method = "norm.predict" [mice].
@@ 2.3.2 Modifications for Multiple Imputation
When using multiple imputation, the imputation models need to be modified one step further. Due to the stochastic nature of multiple imputation, which incorporates variability about the imputation procedure itself, the analysis model outcomes $Y_{j}$ must be included in the imputation models. Otherwise, the imputation models will not be congenial with the analysis model, and the resulting multiple imputation estimates will be biased [DAgostinoMcGowan2024, Moons2006].
There are various ways one might incorporate randomness into the multiple rounds of imputation. One of the earliest and most common methods is to (i) resample the coefficients $\widetilde{\boldsymbol{\alpha}}$ from a distribution based on the fitted ones $\widehat{\boldsymbol{\alpha}}$ in the validation study and (ii) generate the imputed values $\widehat{X}_{ji}$ from the distribution dictated by $\widetilde{\boldsymbol{\alpha}}$ [rubin2004multiple]. For example, if $X_{ji}$ is continuous, then in each imputation the parameters $\widetilde{\boldsymbol{\alpha}}$ and $\tilde{\sigma}$ will first be randomly drawn from a Normal distribution with mean $\widehat{\boldsymbol{\alpha}}$ and standard deviation $\hat{\sigma}$. Then, patients without $X_{ji}$ will have their imputed values $\widehat{X}_{ji}$ simulated from a Normal distribution with mean based on the imputation model (but evaluated at $\widetilde{\boldsymbol{\alpha}}$) and standard deviation $\tilde{\sigma}$. This process is repeated many times, the analysis model is fit to each imputed dataset, and the estimates from each fitted model are ultimately pooled using Rubin’s rules [rubin2004multiple].
We can again use the mice R package for implementation. Specifically, the mice function can be used with argument method = "norm", and it will include the outcome $Y_{j}$ in the imputation model by default. After multiple rounds of imputation, only two more lines of code are needed to fit the analysis model to each and pool the resulting estimates.
@@ 3.1 Data Generating Mechanism
We generated data for samples of $N=1000$ patients, $n$ of whom had their exposures validated. Each patient had a binary confounder $Z$ generated from a Bernoulli distribution with $\Pr(Z=1)=0.3$. Then, true exposures $\boldsymbol{X}=(X_{1},X_{2},X_{3},X_{4},X_{5})$ were generated from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and one of three covariance matrices: independence: $\boldsymbol{\Sigma}_{X}=\boldsymbol{I}_{5}$, equal dependence: $\boldsymbol{\Sigma}_{X}=0.5\boldsymbol{I}_{5}+0.5\boldsymbol{J}_{5}$, and unequal dependence:
@@
where $\boldsymbol{I}_{5}$ the $5\times 5$ identity matrix and $\boldsymbol{J}_{5}$ is a $5\times 5$ matrix with $1$s on the off-diagonal and 0s on the diagonal. Each exposure in $\boldsymbol{X}$ was related to its own numeric response:
@@
for a vector of independent random errors $\boldsymbol{\epsilon}$ generated from a standard multivariate normal distribution. The error-prone versions of these exposures were constructed as $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$, where the measurement errors $\boldsymbol{U}$ were generated from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and covariance matrix $\sigma_{U}^{2}\boldsymbol{I}_{5}$ for error variances $\sigma_{U}^{2}\in\{0.06,0.25,1\}$.
The three validation study designs from Section 2.2 (SRS, ETS-$X_{1}^{*}$, and ETS-$PC_{1}^{*}$) were considered to select $n$ patients to have non-missing $\boldsymbol{X}$ values (i.e., be validated). We focus on comparing the empirical efficiency for the coefficient estimates $\widehat{{\boldsymbol{\beta}}}_{1}$ corresponding to the exposures $\boldsymbol{X}$ under these designs. Specifically, for each of the $1000$ simulated datasets, separate mean models were estimated following the single imputation approach from Section 2.3. These replicates are summarized by the empirical efficiencies for $\widehat{\beta}_{1j}$ ($j\in\{1,2,3,4,5\}$) in each of the five models (i.e., the inverse of the empirical variances of the coefficient estimates).
@@ 3.2 Varied Exposures’ Covariance Structure
The imputation estimator was empirically unbiased for all models, designs, and covariance structures (Web Appendix Figure S1). As expected, ETS-$PC_{1}^{*}$ was most efficient when the exposures $\boldsymbol{X}^{*}$ were correlated (Figure 1A). Under equal dependence, it had similar, large efficiency gains when modeling outcomes $Y_{2},\dots,Y_{5}$ and only fell slightly behind ETS-$X_{1}^{*}$ for $Y_{1}$. Similar observations held under unequal dependence, except that the gap between ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ was wider for the $Y_{1}$ model. In the other models, the efficiency gains for ETS-$PC_{1}^{*}$ were proportional to the average correlation between $X_{j}$ and the other exposures.
Figure 1: Simulation results comparing the empirical efficiency under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. In A), three different covariance structures for the five covariates $X_{1},\dots,X_{5}$ were considered. In B), three different variances $\sigma_{U}^{2}$ for the additive measurement errors $U_{1},\dots,U_{5}$ in covariates $X_{1},\dots,X_{5}$ were considered. In C), three different proportions of validated patients out of $N=1000$ were considered.
@@ 3.3 Varied Exposures’ Error Severity
To isolate the contribution of the exposure error severity, we varied $\sigma_{U}\in\{0.25,0.5,1\}$. We fixed $n=100$ and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as described in Section 3.1. Again, estimates were empirically unbiased for all models and designs (Web Appendix Figure S2).
As $\sigma_{U}$ increased and the measurement errors in $\boldsymbol{X}^{*}$ worsened, the efficiency of all designs and models decreased (Figure 1B). Intuitively, the need for validation is greater when $\sigma_{U}$ is larger, such that $\boldsymbol{X}^{*}$ is “noisier” about $\boldsymbol{X}$. Meanwhile, the benefits of our proposed ETS-$PC_{1}^{*}$ design grew as $\sigma_{U}$ increased, offering pronounced efficiency gains in nearly all models.
@@ 3.4 Varied Validation Proportion
We varied the proportion of the $N=1000$ patients whose exposures $\boldsymbol{X}$ were validated between 10% (used elsewhere), 15%, and 25%. We kept $\sigma_{U}=1$ and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as described in Section 3.1. The single imputation estimates were empirically unbiased for all models with even just 10% validated patients (Web Appendix Figure S3), regardless of the validation study design.
As expected, the efficiency of all designs increased as more patients were validated and thus less data were missing (Figure 1C). The gains with the ETS-$PC_{1}^{*}$ design over ETS-$X_{1}^{*}$ and SRS were persistent and large for all validation proportions considered. Thus, regardless of how many patients are being validated, the ETS-$PC_{1}^{*}$ design gained more information across all models, offering better statistical efficiency, than the ETS-$X_{1}^{*}$ and SRS.
@@ 3.5 Shared Outcome Model
We considered an alternative data generating mechanism where the five exposures $X_{1},\dots,X_{5}$ shared a common numeric outcome $Y$. Under this mechanism, the shared outcome was generated as $Y=\beta_{1}X_{1}+\beta_{2}X_{2}+\beta_{3}X_{3}+\beta_{4}X_{4}+\beta_{5}X_{5}+0.1Z+\epsilon$, where random errors $\epsilon$ come from a standard normal distribution. We fixed $n=100$, $\sigma_{U}=1$, and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as in Section 3.1. Three different scenarios were considered: (i) only $X_{1}$ has a relationship with $Y$ ($\beta_{1}=1$ and all other $\beta_{j}=0$), (ii) only $X_{2}$ has a relationship with $Y$ ($\beta_{2}=1$ and all other $\beta_{j}=0$), and (iii) all $X_{j}$ have relationships with $Y$ (all $\beta_{j}=0.2$). In all scenarios, separate linear models of $Y$ on $X_{j}$ and $Z$ ($j\in\{1,2,3,4,5\}$) were fit using single imputation. We compared the validation study designs when only one exposure was associated with $Y$ (and when it was/was not the one targeted with the ETS-$X_{1}^{*}$ design) versus when all exposures were associated.
Single imputation continued to lead to unbiased estimates with a shared outcome $Y$ (Web Appendix Figure S4). An interesting story emerges when comparing the efficiency of the different validation study designs. Essentially, ETS-$PC_{1}^{*}$ was the “safest” choice, since it offered the same if not better efficiency as ETS-$X_{1}^{*}$ for all models (Figure 2). As expected, when only $X_{1}$ was associated with $Y$, the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs offered nearly identical efficiency in the model using $X_{1}$. However, the ETS-$PC_{1}^{*}$ design offered noticeable efficiency gains over ETS-$X_{1}^{*}$ for the other models. When only $X_{2}$ was associated with $Y$, the ETS-$PC_{1}^{*}$ design offered substantially better efficiency for all five models, but especially for the model with $X_{2}$. Finally, when all five exposures were associated with $Y$, the ETS-$PC_{1}^{*}$ design outperformed the ETS-$X_{1}^{*}$ and SRS designs for all models.
Figure 2: Simulation results comparing the empirical efficiency under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. There was a shared outcome $Y$, and it was generated from the covariates $X_{1},\dots,X_{5}$ under scenarios where only one covariate is associated (only $\beta_{1}\neq 0$ or only $\beta_{2}\neq 0$) versus all covariates are associated (all $\beta_{j}\neq 0$).

### v2

@@ 2.4 Imputation Procedure
In our validated subset, we can estimate the $j$th linear regression model in (1) using the observed exposure $X_{ji}$ directly. For the rest of the patients in the study, we will instead replace their missing value with an informative placeholder $\widehat{X}_{ji}$. Specifically, we adopt a single deterministic imputation framework in the Simulations (Section 3). (Single imputation is computationally quicker than multiple, and we focused on the estimates’ empirical variability to compare the designs.) However, multiple imputation is preferred if estimated variances are desired (e.g., for inference), in which case the imputation models would need to be modified. R code for imputation can be found online at https://github.com/sarahlotspeich/ETS_PCA.
@@ 2.4.1 Different Imputation Models Under Different Validation Study Designs
For an unvalidated patient, $\widehat{X}_{ji}$ comes from an imputation model fit to the validation study. Any fully available information that helps estimate $X_{ji}$ can be incorporated, and any variables used in the validation study design must be included [46].
Under an SRS design, the imputed values are defined simply from the error-prone exposure $X_{ji}^{*}$ and the additional confounders ${\mathbf{Z}}_{i}$ as
@@
When $X_{ji}$ is numeric, the estimated coefficients in (2) can be obtained via normal linear regression using only data from the validation subset. Of course, different regression models could instead be used to impute other types of exposures.
If the validation study was selected under an ETS-$X_{j}^{*}$ design, the error-prone exposure $X_{j}^{*}$ used in the design must be included in all exposures’ imputation models. Let $j^{\prime}$ index which model was used in the design ($j^{\prime}\in\{1,\dots,J\}$). Then, the imputed values $\widehat{X}_{ji}$ are:
@@
Note that the coefficients will take on different values in the first versus second case.
@@
$\widehat{X}_{ji}$ $=\widehat{\alpha}_{0j}+\widehat{\alpha}_{1j}X^{*}_{ji}+\widehat{\alpha}_{2j}Z_{i}+\widehat{\alpha}_{3j}PC^{*}_{1i}.$ (4)
As with the SRS design, estimating the imputation models under either ETS design depends on the type of covariates, but can be done using standard regression techniques. For ease of notation, $\widehat{\boldsymbol{\alpha}}$ are the estimated coefficients for any of the imputation models.
Single imputation following (2)–(4) has been called deterministic imputation, since missing values are replaced with a single model prediction $\widehat{X}_{ji}$ and treated as fixed in the subsequent analyses [6]. This approach can be implemented using existing software to (i) fit the imputation model and (ii) calculate the predictions. Alternatively, the mice function in the mice R package can be used with argument method = "norm.predict" [41].
@@ 2.4.2 Modifications for Multiple Imputation
When using multiple imputation, the imputation models need to be modified one step further. Due to the stochastic nature of multiple imputation, which incorporates variability about the imputation procedure itself, the analysis model outcomes $Y_{j}$ must be included in the imputation models. Otherwise, the imputation models will not be congenial with the analysis model, and the resulting multiple imputation estimates will be biased [6, 26].
There are various ways one might incorporate randomness into the multiple rounds of imputation. One of the earliest and most common methods is to (i) resample the coefficients $\widetilde{\boldsymbol{\alpha}}$ from a distribution based on the fitted ones $\widehat{\boldsymbol{\alpha}}$ in the validation study and (ii) generate the imputed values $\widehat{X}_{ji}$ from the distribution dictated by $\widetilde{\boldsymbol{\alpha}}$ [30]. For example, if $X_{ji}$ is continuous, then in each imputation the parameters $\widetilde{\boldsymbol{\alpha}}$ and $\tilde{\Sigma}$ will first be randomly drawn from a multivariate Normal distribution with mean vector $\widehat{\boldsymbol{\alpha}}$ and variance-covariance matrix $\widehat{\Sigma}$. This process is repeated many times, the analysis model is fit to each imputed dataset, and the estimates from each fitted model are ultimately pooled using Rubin’s rules [30]. We can again use the mice R package. Specifically, the mice function can be used with argument method = "norm", and it will include the outcome $Y_{j}$ in the imputation model by default.
@@ 3.1 Data Generating Mechanism
We generated data for samples of $N=1000$ patients, $n$ of whom had their exposures validated ($n\in\{100,150,250\}$). Each patient had a binary confounder $Z$ generated from a Bernoulli distribution with $\Pr(Z=1)=0.3$. Then, true exposures $\boldsymbol{X}=(X_{1},X_{2},X_{3},X_{4},X_{5})$ were generated from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and one of three covariance matrices: independence: $\boldsymbol{\Sigma}_{X}=\boldsymbol{I}_{5}$, equal dependence: $\boldsymbol{\Sigma}_{X}=0.5\boldsymbol{I}_{5}+0.5\boldsymbol{J}_{5}$, or unequal dependence:
@@
where $\mathbf{0}_{5}$ is a $5\times 1$ vector of zeros, $\boldsymbol{I}_{5}$ the $5\times 5$ identity matrix and $\boldsymbol{J}_{5}$ is a $5\times 5$ matrix with ones on the off-diagonal and zeros on the diagonal. Each exposure in $\boldsymbol{X}$ was related to its own numeric response:
@@
for a vector of independent random errors $\boldsymbol{\epsilon}$ generated from a standard multivariate normal distribution. The error-prone versions of these exposures were constructed as $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$, where the measurement errors $\boldsymbol{U}$ were generated from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and covariance matrix $\sigma_{U}^{2}\boldsymbol{I}_{5}$ ($\sigma_{U}^{2}\in\{0.06,0.25,1\}$).
3.2 Metrics for Comparison
The SRS, ETS-$X_{1}^{*}$, and ETS-$PC_{1}^{*}$ validation study designs were considered. Without loss of generality, the model of $Y_{1}$ given $(X_{1},Z)$ was treated as primary. When discussing individual models, we focus on comparing the empirical efficiency for the coefficient estimates $\widehat{{\boldsymbol{\beta}}}_{1}$ corresponding to the exposures $\boldsymbol{X}$. When evaluating design performance across all modeling objectives simultaneously, we examine the sum of these coefficients’ empirical variances.
For each of $1000$ simulated datasets per setting, separate linear regression models were estimated following the single imputation approach from Section 2.4. These replicates are summarized by the empirical efficiencies for $\widehat{\beta}_{1j}$ ($j\in\{1,2,3,4,5\}$) in each of the five models (i.e., the inverse of the empirical variances) and the empirical total coefficient variability across all models, $\sum_{j=1}^{5}\widehat{\textrm{V}}(\hat{\beta}_{1j})$.
@@ 3.3 Varied Exposures’ Covariance Structure
The imputation estimator was empirically unbiased for all models, designs, and covariance structures (Supplemental Figure S1). As expected, the total empirical variability across all models was smallest for the ETS-$PC_{1}^{*}$ design when the exposures $\boldsymbol{X}$ were dependent (either equally or unequally), such that $PC_{1}^{*}$ was informative about $\boldsymbol{X}^{*}$ and, to a lesser extent, $\boldsymbol{X}$ (Figure 3A). Only when the exposures were independent did the ETS-$X_{1}^{*}$ design lead to a slightly smaller total, and ETS-$PC_{1}^{*}$ still beat SRS in that case.
Figure 3: Simulation results comparing the empirical total coefficient variability across all models $\sum_{j=1}^{5}\widehat{\textrm{V}}(\hat{\beta}_{1j})$ under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. In A), three different covariance structures for the five exposures $X_{1},\dots,X_{5}$ were considered. In B), three different variances $\sigma_{U}^{2}$ for the additive measurement errors $U_{1},\dots,U_{5}$ in exposures $X_{1},\dots,X_{5}$ were considered. In C), three different proportions of validated patients out of $N=1000$ were considered.
When broken down by model, there were clear efficiency gains for all models (of $Y_{2},\dots,Y_{5}$) with the ETS-$PC_{1}^{*}$ design (Figure 4A). These gains were comparable in size across all secondary models under equal dependence, and there was hardly any efficiency loss to the ETS-$X_{1}^{*}$ for the primary model. Under unequal dependence, the largest efficiency gain was seen for the model of $Y_{5}$, which had the exposure $X_{5}$ with the highest correlations, on average, to the others. Under unequal covariance, the efficiency for the primary analysis when validating with ETS-$PC_{1}^{*}$ fell farther behind the ETS-$X_{1}^{*}$ design. However, this model-specific loss was offset by lower total empirical variability across all models, as desired. Only when the exposures were independent did the ETS-$PC_{1}^{*}$ design lose to ETS-$X_{1}^{*}$ at both the model-specific and all-model levels, and it still beat SRS.
Figure 4: Simulation results comparing the empirical efficiency under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. In A), three different covariance structures for the five covariates $X_{1},\dots,X_{5}$ were considered. In B), three different variances $\sigma_{U}^{2}$ for the additive measurement errors $U_{1},\dots,U_{5}$ in exposures $X_{1},\dots,X_{5}$ were considered. In C), three different proportions of validated patients out of $N=1000$ were considered.
@@ 3.4 Varied Exposures’ Error Severity
To isolate the contribution of the exposure error severity, we changed the variance $\sigma_{U}^{2}\in\{0.25,0.5,1\}$ used to generate the errors $\boldsymbol{U}$ and create error-prone exposures $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$. We fixed the validation study size at $n=100$ and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as described in Section 3.1. Again, all estimates were empirically unbiased (Supplemental Figure S2).
The ETS-$PC_{1}^{*}$ offered the smallest total empirical coefficient variability across all models regardless of how severe the measurement error was (Figure 3B). The efficiency gap between the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs widened as $\sigma_{U}^{2}$ increased. Intuitively, the need for validation is greater when $\sigma^{2}_{U}$ is larger, such that $\boldsymbol{X}^{*}$ is noisier about $\boldsymbol{X}$. Meanwhile, the benefits of our proposed ETS-$PC_{1}^{*}$ design over ETS-$X_{1}^{*}$ and SRS grew as $\sigma_{U}^{2}$ increased. In other words, including information from all error-prone exposures, rather than just the primary one, seemed to become more important as the errors worsened.
The same trend was seen when breaking down the efficiency comparison by individual models, as well. As $\sigma_{U}^{2}$ increased, the efficiency of all designs and models decreased, and the advantages of the ETS-$PC_{1}^{*}$ design became more pronounced (Figure 4B). In fact, the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs offered almost the same efficiency in the primary model under the most severe setting where $\sigma_{U}^{2}=1$.
@@ 3.5 Varied Validation Proportion
We varied the proportion of the $N=1000$ patients whose exposures $\boldsymbol{X}$ were validated between $10\%$ (used elsewhere), $15\%$, and 25%. We kept $\sigma_{U}^{2}=1$ and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as described in Section 3.1. The estimates were empirically unbiased with even just $10\%$ validated (Supplemental Figure S3).
For all validation proportions considered, the ETS-$PC_{1}^{*}$ design offered the smallest empirical total variability across the five models (Figure 3C). The gap between the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs was fairly consistent in magnitude across all three settings. As the validation proportion increased, fewer patients have missing data and need imputation, leading to better efficiency for all designs. Still, incorporating all five error-prone exposures through the ETS-$PC_{1}^{*}$ design remained beneficial. Regardless of how many patients were validated, the ETS-$PC_{1}^{*}$ design gained more information across all models, offering better statistical efficiency, than the ETS-$X_{1}^{*}$ and SRS.
When evaluating the models individually, the efficiency of all designs increased as more patients were validated, as expected (Figure 4C). The gains with the ETS-$PC_{1}^{*}$ design over ETS-$X_{1}^{*}$ and SRS were persistent and large for all validation proportions considered. This finding suggests that the ETS-$PC_{1}^{*}$ design could be beneficial for partial validation studies with a wide range of budgets and sizes.
@@ 3.6 Shared Outcome Model
We considered an alternative data generating mechanism where the five exposures $X_{1},\dots,X_{5}$ shared a common numeric outcome $Y$. Under this mechanism, the shared outcome was generated as $Y=\beta_{1}X_{1}+\beta_{2}X_{2}+\beta_{3}X_{3}+\beta_{4}X_{4}+\beta_{5}X_{5}+0.1Z+\epsilon$, where random errors $\epsilon$ come from a standard normal distribution. We fixed the validation study size $n=100$, error variance $\sigma_{U}^{2}=1$, and used the equal dependence covariance structure for $\boldsymbol{X}$; all other variables were generated as in Section 3.1. Three scenarios were included: (i) only $X_{1}$ has a relationship with $Y$ ($\beta_{1}=1$ and all other $\beta_{j}=0$), (ii) only $X_{2}$ has a relationship with $Y$ ($\beta_{2}=1$ and all other $\beta_{j}=0$), and (iii) all $\boldsymbol{X}$ have relationships with $Y$ (all $\beta_{1}=\cdots=\beta_{5}=0.2$). In all scenarios, separate linear models of $Y$ on $X_{j}$ and $Z$ were fit using single imputation. Through these settings, we compared the validation study designs when only one exposure was associated with $Y$ (and when it was/was not the one targeted with the ETS-$X_{1}^{*}$ design) versus when all exposures were associated.
Imputation continued to lead to unbiased estimates (Supplemental Figure S4), and an interesting new story emerges when comparing the designs’ efficiency. Essentially, ETS-$PC_{1}^{*}$ was the “safest” choice, since it offered the same if not better efficiency as ETS-$X_{1}^{*}$, both in terms of the total across all coefficients and in the individual models (Figure 5). As expected, when only $X_{1}$ was associated with $Y$, the ETS-$PC_{1}^{*}$ and ETS-$X_{1}^{*}$ designs offered nearly identical efficiency in the model using $X_{1}$. However, the ETS-$PC_{1}^{*}$ design offered noticeable efficiency gains over ETS-$X_{1}^{*}$ for the other models. When only $X_{2}$ was associated with $Y$, the ETS-$PC_{1}^{*}$ design offered substantially better efficiency in total and for all five models separately, but especially for the model with $X_{2}$. Finally, when all $\boldsymbol{X}$ were associated with $Y$, the ETS-$PC_{1}^{*}$ design outperformed the ETS-$X_{1}^{*}$ and SRS designs for all models.
Figure 5: Simulation results comparing the empirical total coefficient variability across all models (A) and empirical efficiency for the exposure coefficient per model (B) under simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on the first principal component (ETS-$PC_{1}^{*}$) validation study designs. There was a shared outcome $Y$, and it was generated from the exposures $X_{1},\dots,X_{5}$ under scenarios where only one exposure was associated (only $\beta_{1}\neq 0$ for $X_{1}$ or only $\beta_{2}\neq 0$ for $X_{2}$) versus all covariates are associated (all $\beta_{j}\neq 0$ for $X_{1},\dots,X_{5}$).